# VRAIL: Vectorized Reward-based Attribution for Interpretable Learning

### v3

@@ 1 Introduction
A toy experiment shown in fig. 1 further illustrates the necessity of our proposed method. In the standard Taxi-v3 environment from the OpenAI Gymnasium library Towers et al. [2024], the agent’s observations do not include explicit wall information, which plays a critical role in successful navigation. As shown in fig. 1, without this structural information, a vanilla DQN baseline fails to converge in 2 out of 10 random seeds, highlighting the instability that can arise from incomplete state representations. Although we found that manually adding wall information resolved the issue and led to consistent convergence, such interventions are impractical in real-world settings. Our expectation is that VRAIL achieves stable performance even in the original, unmodified environment by learning to attribute value to latent features within state transitions. Therefore, VRAIL will offer a practical and potentially generalizable approach for RL tasks where full state observability cannot be guaranteed.
@@ 2 Method
We propose VRAIL (Vectorized Reward-based Attribution for Interpretable Learning), a framework for learning interpretable weight via bi-level optimization. As illustrated in fig. 2, the process consists of two interacting stages: a reinforcement learning (RL) stage and a deep learning (DL) stage. The overall objective is to help the agent recognize meaningful feature transitions as implicit subgoals, improving sample efficiency and interpretability.
@@ 3.2 Training Stability and Convergence Speed
We evaluate the convergence behavior of VRAIL in terms of both training stability and speed, using a standard DQN baseline for comparison. A run is considered converged if the agent consistently achieves a moving average reward above a fixed threshold during training. As shown in fig. 3, the baseline DQN converged in 8 out of 10 runs. In contrast, Linear VRAIL converged in all 10 runs and Quadratic VRAIL converged in 9 out of 10 runs, demonstrating greater robustness than the baseline.
@@
We evaluate convergence speed by measuring the number of training epochs required to reach moving average reward thresholds of $-10$, $-5$, 00, and $+5$. As shown in table 1, Quadratic VRAIL consistently reaches all thresholds faster than both DQN and Linear VRAIL. While Linear VRAIL converges more reliably, it takes slightly longer than DQN to achieve the same rewards. These results highlight a trade-off: Linear VRAIL offers robustness, while Quadratic VRAIL combines speed with stability.
@@ 3.3 Effect of Shaped Reward Transfer
To further isolate the effect of learned reward shaping, we evaluate whether the reward function learned by VRAIL can benefit standard RL agents when applied independently. Specifically, we pretrain Linear VRAIL to learn a potential-based reward function, then fix this function and apply it to a vanilla DQN agent without further reward updates. As shown in fig. 4, the baseline DQN fails to converge in 1 out of 5 seeds. In contrast, DQN trained with the reward shaped by pretrained Linear VRAIL successfully converges in all 5 seeds. This suggests that VRAIL’s learned reward function serves as a robust shaping mechanism, improving convergence stability even when decoupled from the full bi-level optimization loop.
@@ 4.1 Linear VRAIL
In Linear VRAIL, the value function is represented as $\hat{V}(x_{s};w)=w^{\top}x_{s}$, where $w\in\mathbb{R}^{19}$ is a weight vector, with each element corresponding to a state feature. To capture a range of learned behaviors across different runs, we average the learned weight vectors over 10 random seeds and visualize the resulting feature importance. As shown in fig. 5(a), the feature assigned the highest weight corresponds to whether the passenger is currently in the taxi, represented as "P: T". This suggests that the model implicitly identifies passenger possession as a key subgoal. Notably, this aligns with the structure of the Taxi environment, where incorrect pickup or drop-off actions incur a penalty of $-10$, thereby reinforcing proper pickup and drop-off.
@@ 4.2 Quadratic VRAIL
In Quadratic VRAIL, the value function is defined as $\hat{V}(x_{s};W)=x_{s}^{\top}Wx_{s}$, where $W\in\mathbb{R}^{19\times 19}$ captures pairwise interactions between state features. We averaged the learned $W$ over 9 converged seeds and visualized it as a heatmap. In fig. 5(b), each axis represents a state feature, and cell intensity reflects the strength of interaction between feature pairs.
@@ Appendix A Team member Contribution
  • Jina Kim : Backbone model implementation (Linear VRAIL), DQN toy experiment (Fig 1), self-attention based model trials.
@@ Appendix C Trial and Error
  • Environment variants: Sparse reward version (Only +20 when arrival, $-10$ for illegal actions) was tested, but didn’t work well at all.

### v4

@@ 1 Introduction
A toy experiment shown in fig.˜1 further illustrates the necessity of our proposed method. In the standard Taxi-v3 environment from the OpenAI Gymnasium library Towers et al. [2024], the agent’s observations do not include explicit wall information, which plays a critical role in successful navigation. As shown in fig.˜1, without this structural information, a vanilla DQN baseline fails to converge in 2 out of 10 random seeds, highlighting the instability that can arise from incomplete state representations. Although we found that manually adding wall information resolved the issue and led to consistent convergence, such interventions are impractical in real-world settings. Our expectation is that VRAIL achieves stable performance even in the original, unmodified environment by learning to attribute value to latent features within state transitions. Therefore, VRAIL will offer a practical and potentially generalizable approach for RL tasks where full state observability cannot be guaranteed.
@@ 2 Method
We propose VRAIL (Vectorized Reward-based Attribution for Interpretable Learning), a framework for learning interpretable weight via bi-level optimization. As illustrated in fig.˜2, the process consists of two interacting stages: a reinforcement learning (RL) stage and a deep learning (DL) stage. The overall objective is to help the agent recognize meaningful feature transitions as implicit subgoals, improving sample efficiency and interpretability.
@@ 3.2 Training Stability and Convergence Speed
We evaluate the convergence behavior of VRAIL in terms of both training stability and speed, using a standard DQN baseline for comparison. A run is considered converged if the agent consistently achieves a moving average reward above a fixed threshold during training. As shown in fig.˜3, the baseline DQN converged in 8 out of 10 runs. In contrast, Linear VRAIL converged in all 10 runs and Quadratic VRAIL converged in 9 out of 10 runs, demonstrating greater robustness than the baseline.
@@
We evaluate convergence speed by measuring the number of training epochs required to reach moving average reward thresholds of 10-10, 5-5, 0, and +5+5. As shown in table˜1, Quadratic VRAIL consistently reaches all thresholds faster than both DQN and Linear VRAIL. While Linear VRAIL converges more reliably, it takes slightly longer than DQN to achieve the same rewards. These results highlight a trade-off: Linear VRAIL offers robustness, while Quadratic VRAIL combines speed with stability.
@@ 3.3 Effect of Shaped Reward Transfer
To further isolate the effect of learned reward shaping, we evaluate whether the reward function learned by VRAIL can benefit standard RL agents when applied independently. Specifically, we pretrain Linear VRAIL to learn a potential-based reward function, then fix this function and apply it to a vanilla DQN agent without further reward updates. As shown in fig.˜4, the baseline DQN fails to converge in 1 out of 5 seeds. In contrast, DQN trained with the reward shaped by pretrained Linear VRAIL successfully converges in all 5 seeds. This suggests that VRAIL’s learned reward function serves as a robust shaping mechanism, improving convergence stability even when decoupled from the full bi-level optimization loop.
@@ 4.1 Linear VRAIL
In Linear VRAIL, the value function is represented as $\hat{V}(x_{s};w)=w^{\top}x_{s}$, where $w\in\mathbb{R}^{19}$ is a weight vector, with each element corresponding to a state feature. To capture a range of learned behaviors across different runs, we average the learned weight vectors over 10 random seeds and visualize the resulting feature importance. As shown in fig.˜5(a), the feature assigned the highest weight corresponds to whether the passenger is currently in the taxi, represented as "P: T". This suggests that the model implicitly identifies passenger possession as a key subgoal. Notably, this aligns with the structure of the Taxi environment, where incorrect pickup or drop-off actions incur a penalty of 10-10, thereby reinforcing proper pickup and drop-off.
@@ 4.2 Quadratic VRAIL
In Quadratic VRAIL, the value function is defined as $\hat{V}(x_{s};W)=x_{s}^{\top}Wx_{s}$, where $W\in\mathbb{R}^{19\times 19}$ captures pairwise interactions between state features. We averaged the learned $W$ over 9 converged seeds and visualized it as a heatmap. In fig.˜5(b), each axis represents a state feature, and cell intensity reflects the strength of interaction between feature pairs.
@@ Appendix A Team member Contribution
  • Jina Kim : Backbone model implementation (Linear VRAIL), DQN toy experiment (Fig 1), self-attention based model trials, final refinement of the report.
@@ Appendix C Trial and Error
  • Environment variants: Sparse reward version (Only +20 when arrival, 10-10 for illegal actions) was tested, but didn’t work well at all.